Circular No. 6763

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997dn IN NGC 3451 G. M. Hurst, Basingstoke, England, reports the discovery of an appar-

ent supernova (mag ∼ 16.0) by Tom Boles, Wellingborough, on CCD frames taken during Oct. 29.177–29.217 UT as part of the U.K. Nova/Supernova Patrol. The star is also present at a similar brightness on an image taken by Boles on Oct. 30.177. N. D. James, Chelmsford, provides the following astrometry for SN 1997dn from Boles’ discovery image: α = 10^{h}54^{m}18^{s}.95, δ = +27^{o}14′07′′.3 (equinox 2000.0), which is ∼ 35′′ west and 20′′ south of the center of NGC 3451. No star is present at this location on a master CCD image of NGC 3451 taken on 1997 May 31 by Boles.

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum of SN 1997dn taken by P. Berlind at the Whipple Observatory 1.5-m telescope on Oct. 30.5 UT indicates that it is a type-II supernova after maximum. A prominent P-Cyg profile of Hα is visible, with a photospheric expansion velocity (measured from peak to trough) of 6500 km/s. Also present in the spectrum are Hβ, Hγ, He I (587.5 nm), and Fe II. The redshift of the host galaxy is 1300 km/s, measured from narrow Hα emission.

CATACLYSMIC VARIABLE IN CETUS The Beijing Astronomical Observatory (BAO) supernova survey, with

A. Esamdin, Urumqi Astronomical Station, reports their discovery on Oct. 28, using the 0.6-m reflector at Xinglong Station, of a cataclysmic variable star in outburst at α = 0^{h}27^{m}27^{s}.93, δ = −1^{o}08′26′′.8 (equinox 2000.0). Available R magnitudes: Oct. 17.64 UT, 14.5; 21.67, 14.9; 28.61, 15.4; 29.61, 15.6; 30.45, 15.8. The variable did not apppear on CCD images obtained on Oct. 8 (limiting mag 18.5), and it appears as a very faint object on the original Palomar Sky Survey. A spectrum of the star, obtained by X. Dong and Y. Zhao with the BAO 2.16-m telescope at Xinglong on Oct. 30.56, shows typical features of cataclysmic variables in outburst, with mainly a blue continnum at 500–800 nm, and a spectrum at 380–500 nm that is similar to the dwarf nova CY Lyr at maximum, with absorption of Hδ, Hγ, Hβ, and weak He I (447.1 nm).

COMET C/1997 D1 (MUELLER) Visual m_{1} estimates by K. Hornoch, Lelekovice, Czech Republic (0.35-

m reflector): Sept. 29.12 UT, 12.7; Oct. 6.11, 12.3; 28.03, 11.9; 28.97, 11.8.

1997 October 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green